# Attention is All They Need: Exploring the Media Archaeology of the Computer Vision Research Paper

### v2

@@ Abstract.
Research papers, in addition to textual documents, are a designed interface through which researchers communicate. Recently, rapid growth has transformed that interface in many fields of computing. In this work, we examine the effects of this growth from a media archaeology perspective, through the changes to figures and tables in research papers. Specifically, we study these changes in computer vision over the past decade, as the deep learning revolution has driven unprecedented growth in the discipline. We ground our investigation through interviews with veteran researchers spanning computer vision, graphics and visualization. Our analysis focuses on the research attention economy: how research paper elements contribute towards advertising, measuring and disseminating an increasingly commodified “contribution.” Through this work, we seek to motivate future discussion surrounding the design of both the research paper itself as well as the larger sociotechnical research publishing system, including tools for finding, reading and writing research papers.
@@ 1. Introduction
By attention, we evoke the attention economy, first proposed by Herbert Simon (Simon, 1971) and applied to the internet by authors such as Goldhaber (Goldhaber, 1997) and Davenport and Beck (Davenport and Beck, 2001). An attention economy in the workplace has been discussed in CSCW, for example in the work of Yardi et al. on the attention economy of internal corporate blogs (Yardi et al., 2009) or Kraut et al. on the attention economy of email marketing (Kraut et al., 2002). To understand its commodification, we turn to the post-Marxist analysis of Claudio Celis Bueno (Bueno, 2017). Bueno frames online attention, not just as a scarce resource, but as a form of labor—human activity which has the potential to generate surplus value. We claim that this theory of attention as labor applies to the attention paid to academic research as well, and helps to explain recent changes to computer vision.
@@
Perhaps the biggest hint that the design of computer vision papers has changed with the rapid changes in its disciplinary culture is scholarly humor. In the 2010 computer vision satire paper “Paper Gestalt” (Von Bearnensquash, 2010), the pseudonymous authors “take the simple intuition that the quality of a paper can be estimated by merely glancing through the general layout, and use this intuition to build a system that employs basic computer vision techniques to predict if the paper should be accepted or rejected,” (Von Bearnensquash, 2010) and suggest that this system might replace the peer review process. The visual design of the paper parodies the qualities it observes, including unnecessary complex equations and long algorithms.
@@
In the following sections, we examine historical computer vision research papers from a media archaeology perspective, focused on the way their visual style has developed over time. To better understand these developments in context, we also report results of interviews with twelve veteran researchers whose work spans computer vision, graphics, and visualization. Our approach focuses on elements of the research paper and larger publication system based on their roles in advertising, measuring, and disseminating an increasingly commodified “contribution.” We find that as academic discourse has moved online, the limiting factor on publishing has shifted from printing costs to the attention of other researchers, which has changed both the culture of computer vision, as well as the design of its research papers. While this finding echoes existing discourse surrounding “publish or perish” (Garfield, 1996; Lee, 2014), the “marketization” of academic discourse (Fairclough, 1993), use of “hype” in science communication (Soto-Sanfiel et al., 2023) or promotional qualities in titles and abstracts (Hyland, 2023), we emphasize the novelty of our approach. To the best of our knowledge ours is the first connection between these discussions of scholarly discourse and the designed visual artifact of the research paper.
@@ 2.3. The research paper as a site for interaction design research
This research community is also interested in designing for the writing process: Head et al. study ways of augmenting digital documents with definitions of terms and symbols to improve readability (Head et al., 2021), and how authors improve readability by augmenting the visual design of their equations (Head et al., 2022). Manzoor et al. develop a LaTeX editor extension to improve the accessibility of LaTeX for authors with visual impairments (Manzoor et al., 2018) and Hara et al. develop a system for generating Braille documents from mathematical expressions written in LaTeX (Hara et al., 2000). Gobert and Beaudouin-Lafon conduct a study of LaTeX users and design an extension for VSCode that uses transitional representations of document objects like tables to improve the editing experience (Gobert and Beaudouin-Lafon, 2022). Haber et al. study how groups of coworkers interact differently when using physical vs. virtual documents (Haber et al., 2014). While we do not design or develop any technical tools, our work provides additional evidence for the importance of further design studies of research paper reading and writing.
@@ 2.4. The culture of computer vision
Motivated by the dangers of algorithmic discrimination and safety concerns in systems relying on computer vision algorithms, several recent papers have studied the culture of the computer vision research community and its understanding of data and truth, often using research papers as texts for analysis. Su and Crandall study the emotional state of the computer vision community, finding that the deep learning revolution and subsequent growth has had a profound effect on its culture, and leading both to excitement regarding progress as well as to isolation and malaise (Su and Crandall, 2021). Denton et al. use a discourse analysis approach to study the history of the ImageNet dataset through the research papers and presentation slides of Fei-Fei Li (Denton et al., 2021). Scheuerman, Denton, and Hanna study a corpus of 500 papers describing computer vision datasets and analyze the values implicit in their writing: efficiency, universality, impartiality, and model development over dataset development (Scheuerman et al., 2021). While our work also analyzes computer vision though its research papers, our interest is in studying these papers as primarily visual media.
A wide variety of authors have written critically about the culture of data collection and dataset use in machine learning. For example, Sculley et al. and Ethayarajh and Jurafsky critique the concept of benchmarks and leaderboards in machine learning, arguing in different ways that steadily increasing scores do not always correspond to progress (Sculley et al., 2018; Ethayarajh and Jurafsky, 2020). For a more comprehensive survey please see Paullada et al. (Paullada et al., 2021). Our work here, which does not investigate data collection or dataset usage directly, nevertheless echoes these themes.
@@ 3. Methods
Our inquiry began through visual analysis of historical computer vision research papers. We found the changes to the visual design of these papers surprising, particularly the increasing prevalence of highly complex figures, and decided to investigate further by conducting both semi-structured interviews with researchers who had been active in computer vision, graphics, and visualization for several decades, and computational analysis of a corpus of research papers published between 2013 and 2021 in the IEEE Conference on Computer Vision and Pattern Recognition—the highest ranked publication venue in computer science and third highest in all of science, according to Google Scholar (goo, 2024). We study this conference, rather than similar conferences in machine learning or natural language processing, because of the importance of visual presentation in computer vision. Along the way, our visual analysis, interviews and computational findings condensed into a media archaeology approach.
@@ 3.1. Media Archaeology
Inspired by recent work on the history of computer graphics (Gaboury, 2021), we examine historical research papers through the lens of media archaeology. By “archaeology,” this method refers to Foucault’s concept of archaeology as the “uncovering of the archive” (Foucault, 1969, p.131), rather than physical excavation. Media archaeology, as defined by Huhtamo and Parikka, “rummages textual, visual and auditory archives as well as collections of artifacts, emphasizing both the discursive and material manifestations of culture” (Huhtamo and Parikka, 2011, p. 3). In other words, we study digital media objects as components of a media culture; a system of practices and meanings which structure our interpretation (Rose, 2016). We also consider their digital materiality as physical patterns of bits on hard drives, produced through the labor of researchers. By drawing attention to material culture, media archaeology is in dialogue with theories of historical materialism, which consider economic factors, production and consumption of goods and services, as the driving force of history, rather than revolutionary ideas or progress narratives (Gaboury, 2021, p. 5).
@@
In addition to our computational analysis and detailed study of the CVPR proceedings, we “rummaged” research papers from the IEEE, Computer Vision Foundation, and ACM digital archives, including the IEEE International Conference on Computer Vision (ICCV), European Conference on Computer Vision (ECCV), IEEE Transactions on Pattern Analysis and Machine Intelligence (PAMI), and the ACM Speical Interest Group on Graphics and Interactive Techniques conference (SIGGRAPH). We examined ICCV, ECCV, and PAMI in addition to CVPR because they are the four most influential computer vision venues according to Google Scholar,111https://web.archive.org/web/20220709182958/https://scholar.google.com/citations?view_op=top_venues&hl=en&vq=eng_computervisionpatternrecognition and SIGGRAPH due to its historical significance and prevalence in our interviews. We turned to these venues to build context for the results we were seeing in CVPR. While we did not systematically code the enormous number of papers published in these venues, we collected dozens of screenshots of interesting figures and tables and discussed their visual style in weekly meetings, integrating interview and computational results which eventually coalesced into historical narratives.
@@ 3.2. Interviews
To augment our visual analysis, we conducted interviews with veteran computer vision researchers. As there are a relatively small number of eligible participants, we recruited individuals via email and in-person at the CVPR 2022 conference. Participants are listed by years of industry and academic experience in Table 1. This sample is highly non-representative, showing both survival bias based on who remains in the field over a long period of time and selection bias based on who agreed to an interview, so it does not serve as the source of our claims. Instead, it provides important social context to our visual analysis.
@@
After transcribing the interviews, we engaged in a process of iterative memoing, integrated into our weekly meetings where we compared examples from historical research papers and interview excerpts which shared common themes. We focused our analysis on identifying patterns from the context that participants provided while telling stories about their writing, as well how participants thought about their papers in relation to changing disciplinary practices, rather than simply analyzing the papers participants discussed, or discussing our participants’ analysis of their own work.
@@ 3.3. Computational Analysis
As both qualitative approaches rely on the analysis of specific examples, we used supplementary quantitative analysis to verify that phenomena we observed in the interviews and media archaeology were actually as pervasive as participants seemed to think. Specifically, we were interested in whether teaser images, figures and tables were becoming more prevalent in CVPR papers over time, and whether more titles were following a particular format containing an acronym followed by a colon. To answer these questions, we collected open access PDF files from thecvf.com published in CVPR 2013 to 2021 and article metadata from ieeexplore.ieee.org for CVPR 1992 to 2020. We manually inspected PDFs from thecvf.com to count teaser images, then used the Linux pdftotext tool combined with regular expressions to count figures and tables. From the IEEExplore metadata, we used regular expressions to parse paper titles, treating a word with more than two capital letters as an acronym, and an acronym as unique if it only appears once in a given year of data.
@@ 4. Findings
Through our analysis, we found that several elements of the contemporary computer vision research paper serve as material traces of the disciplinary change which took place over the course of the 2010s. We find that papers are written to promote a specific contribution which is increasingly commodified. By “contribution,” we mean the model, algorithm, method, dataset or other system that the paper offers to readers. By commodified, we mean reduced to its exchange value in terms of the attention it generates. As Simon theorized (Simon, 1971), and authors including Goldhaber (Goldhaber, 1997) and Davenport and Beck (Davenport and Beck, 2001) have applied to the web, when “capital, labor, information and knowledge are all in plentiful supply” the limiting factor is the attention of consumers, which begins to be treated like a commodity (Davenport and Beck, 2001).
However, this concept of the attention economy presupposes that attention is valuable, which is unhelpful for understanding why academics seek the attention of their peers instead of the larger research community or general public. To better understand why academic attention is uniquely valuable, we employ Bueno’s explanation, that online attention, in addition to a commodity, is a form of labor (Bueno, 2017). This claim relies on Romano Alquati’s concept of “valorization information” via Pasquinelli (Pasquinelli, 2015). Alquati, in his study of an Olivetti computer factory, found that workers contribute to mostly-automated production processes through their micro-decisions, which served as a feedback-control signal for the production process. Bueno argues that attention can provide a similar signal. Specifically, by monitoring attention, companies can forecast consumer behavior and more efficiently follow market trends. By doing so, these companies extract surplus value from online behavior data (Bueno, 2017, Ch. 2). We believe a similar process may be at play in academic publishing, where authors can leverage highly specialized early attention to increase the prestige value of their research. This process can take the form of feedback, where authors improve their work based on an early online response, or simply where early viewers find a paper interesting and share it, judging it as valuable, putting it in front of more eyes and increasing the chance that it will reach the right audience. Eventually, that attention translates into citations, jobs, students, grants and other benefits of academic prestige.
We study three aspects of contemporary research papers: the teaser image, the results table and the high resolution figure. We have grouped our results into thematic sections surrounding these three elements in relation to the concept of a paper’s contribution.
@@ 4.1.1. Teaser images
The first element of research paper design that we examine is the “teaser image,” a large first-page figure which summarizes the paper. Several examples of teaser images are shown in Figures 1 and 2. These figures are functionally similar to the trend of “visual abstracts” (Ramos and Concepcion, 2020) and “table of contents images” (Buriak, 2011) in the biomedical sciences, however unlike these forms which are graphical summaries of the text, teaser images are part of the main paper and are usually a visualized system output that the authors want to show. Teaser images have been steadily gaining popularity within CVPR, as shown in Figure 3 left. For this analysis, we defined the teaser image as a first-page figure which covers the entire width of the page, not just a column.
@@
P12 ascribes the teaser image to the famous graphics researcher Randy Pausch: “Randy Pausch…takes credit for putting the teaser image in SIGGRAPH papers…he claimed that he did a paper where half the first page was a teaser image and after that that became the norm where people started always putting these images.” The paper in question is pictured in Figure 2 (a). While Pausch did not invent the large first-page figure (Tim Berners-Lee famously used one in his 1989 proposal for the world wide web, for example (Berners-Lee, 1989)), he did publish the first paper in SIGGRAPH with this layout, and it quickly became popular. P2 and P10 point to the quick adoption of teaser images, which even became institutionalized in the SIGGRAPH template, which is now the template for all ACM conferences (Figure 2 (c).
@@
P12 defines the teaser as a visualization of either a result or a system and explains that it has spread to many other conferences because of the way it attracts attention: “It’s a trailer. It’s to get people in…I think it’s a very compelling way to convey what the paper does.” For her, the teaser is a highly effective innovation which improves research papers. P2 echoes that sentiment: “They just made the papers look good! I mean, It’s much more memorable and there are some papers still today that I don’t remember the title, but you see the picture and you’re like, oh yeah, that’s the Randy Pausch paper on the VR for whatever, right? So yeah there’s a few of those that are just like, really iconic first pagers.” P2 is referencing the same Pausch paper as P12 (Pausch et al., 1996), highlighting its memorability.
@@
The teaser image is a trailer, a hook; it advertises the paper to the potential reader. The authors want to promote their paper and showcase the best results they can because the sheer number of and easy access to papers has made it harder to stand out. The visual organization of these figures echoes that theme. Notice the commonalities between the two teaser images in Figure 2 (d) and (e). Both show the output of several algorithms attempting to solve the same problem, but they depict the results in such a way that their method is clearly best. This visual effect mirrors that of advertisements, which depict two competing products in action (Figure 2 (f)). While the experiments in research papers are far more sophisticated than those in advertisements, they share a common visual design. Along these lines, we find a recurring theme of teaser images which rely on the iconography of the 3D-reconstructed, brightly colored human body as a particularly compelling visual. Several examples are shown in Figure 1. Human faces and bodies have been found to attract consumer attention to advertisements (Wilkinson and Light, 2011; Ohme et al., 2011); a similar principle may be motivating computer vision authors.
@@
We observe a similar advertising quality in titles. Figure 3 right shows the rise in popularity of a particular title construction where an acronym, which is usually the name of a model, is followed by a colon: “HOPE-Net: A Graph-Based Model for Hand-Object Pose Estimation” (Doosti et al., 2020) or “DeMoN: Depth and Motion Network for Learning Monocular Stereo” (Ummenhofer et al., 2017). These names both signal that the paper proposes a new model and brands the paper with a short, memorable name which is often a cultural reference or clever pun. An especially cheeky example is Joseph Redmon and Ali Farhadi’s 2017 paper “YOLO9000: Better Faster Stronger” (Redmon and Farhadi, 2017), which improves upon an earlier model called YOLO (short for You Only Look Once) (Redmon et al., 2016), and references a variety of memes from the mid-2010s.222YOLO itself was a meme, short for “you only live once,” the number 9000 references a famous quote from the television show Dragon Ball Z and the phrase “better, faster, stronger” references a Daft Punk song. While unusual in computer vision, this kind of nerdy referential humor has been observed in other areas, for example, web design (Goree et al., 2022).

### v3

@@ Abstract.
Research papers, in addition to textual documents, are a designed interface through which researchers communicate. Recently, rapid growth has transformed that interface in many fields of computing. In this work, we examine the effects of this growth from a media archaeology perspective, through the changes to figures and tables in research papers. Specifically, we study these changes in computer vision over the past decade, as the deep learning revolution has driven unprecedented growth in the discipline. We ground our investigation through interviews with veteran researchers spanning computer vision, graphics, and visualization. Our analysis focuses on the research attention economy: how research paper elements contribute towards advertising, measuring, and disseminating an increasingly commodified “contribution.” Through this work, we seek to motivate future discussion surrounding the design of both the research paper itself as well as the larger sociotechnical research publishing system, including tools for finding, reading, and writing research papers.
@@ 1. Introduction
By attention, we evoke the attention economy, first proposed by Herbert Simon (Simon, 1971) and applied to the internet by authors such as Goldhaber (Goldhaber, 1997) and Davenport and Beck (Davenport and Beck, 2001). An attention economy in the workplace has been discussed in CSCW, for example in the work of Yardi et al. on the attention economy of internal corporate blogs (Yardi et al., 2009) or Kraut et al. on the attention economy of email marketing (Kraut et al., 2002). To understand its commodification, we turn to the post-Marxist analysis of Claudio Celis Bueno (Bueno, 2017). Bueno frames online attention not just as a scarce resource, but as a form of labor—human activity which has the potential to generate surplus value. We claim that this theory of attention as labor applies to the attention paid to academic research as well, and helps to explain recent changes to computer vision.
@@
Perhaps the biggest hint that the design of computer vision papers has changed with the rapid changes in its disciplinary culture is scholarly humor. In the 2010 computer vision satire paper “Paper Gestalt” (Von Bearnensquash, 2010), the pseudonymous authors “take the simple intuition that the quality of a paper can be estimated by merely glancing through the general layout, and use this intuition to build a system that employs basic computer vision techniques to predict if the paper should be accepted or rejected” (Von Bearnensquash, 2010), and suggest that this system might replace the peer review process. The visual design of the paper parodies the qualities it observes, including unnecessary complex equations and long algorithms.
@@
In the following sections, we examine historical computer vision research papers from a media archaeology perspective, focused on the way their visual style has developed over time. To better understand these developments in context, we also report results of interviews with twelve veteran researchers whose work spans computer vision, graphics, and visualization. Our approach focuses on elements of the research paper and larger publication system based on their roles in advertising, measuring, and disseminating an increasingly commodified “contribution.” We find that as academic discourse has moved online, the limiting factor on publishing has shifted from printing costs to the attention of other researchers, which has changed both the culture of computer vision as well as the design of its research papers. While this finding echoes existing discourse surrounding “publish or perish” (Garfield, 1996; Lee, 2014), the “marketization” of academic discourse (Fairclough, 1993), use of “hype” in science communication (Soto-Sanfiel et al., 2023), or promotional qualities in titles and abstracts (Hyland, 2023), we emphasize the novelty of our approach. To the best of our knowledge ours is the first connection between these discussions of scholarly discourse and the designed visual artifact of the research paper.
@@ 2.3. The research paper as a site for interaction design research
This research community is also interested in designing for the writing process: Head et al. study ways of augmenting digital documents with definitions of terms and symbols to improve readability (Head et al., 2021), and how authors improve readability by augmenting the visual design of their equations (Head et al., 2022). Manzoor et al. develop a LaTeX editor extension to improve the accessibility of LaTeX for authors with visual impairments (Manzoor et al., 2018), and Hara et al. develop a system for generating Braille documents from mathematical expressions written in LaTeX (Hara et al., 2000). Gobert and Beaudouin-Lafon conduct a study of LaTeX users and design an extension for VSCode that uses transitional representations of document objects like tables to improve the editing experience (Gobert and Beaudouin-Lafon, 2022). Haber et al. study how groups of coworkers interact differently when using physical vs. virtual documents (Haber et al., 2014). While we do not design or develop any technical tools, our work provides additional evidence for the importance of further design studies of research paper reading and writing.
@@ 2.4. The culture of computer vision
Motivated by the dangers of algorithmic discrimination and safety concerns in systems relying on computer vision algorithms, several recent papers have studied the culture of the computer vision research community and its understanding of data and truth, often using research papers as texts for analysis. Su and Crandall study the emotional state of the computer vision community, finding that the deep learning revolution and subsequent growth have had a profound effect on its culture, leading both to excitement as well as isolation and malaise (Su and Crandall, 2021). Denton et al. use a discourse analysis approach to study the history of the ImageNet dataset through the research papers and presentation slides of Fei-Fei Li (Denton et al., 2021). Scheuerman, Denton, and Hanna study a corpus of 500 papers describing computer vision datasets and analyze the values implicit in their writing: efficiency, universality, impartiality, and model development over dataset development (Scheuerman et al., 2021). While our work also analyzes computer vision though its research papers, our interest is in studying these papers as primarily visual media.
A wide variety of authors have written critically about the culture of data collection and dataset use in machine learning. For example, Sculley et al. and Ethayarajh and Jurafsky critique the concept of benchmarks and leaderboards in machine learning, arguing in different ways that steadily increasing scores do not always correspond to progress (Sculley et al., 2018; Ethayarajh and Jurafsky, 2020). For a more comprehensive survey, please see Paullada et al. (Paullada et al., 2021). Our work here, which does not investigate data collection or dataset usage directly, nevertheless echoes these themes.
@@ 3. Methods
Our inquiry began through visual analysis of historical computer vision research papers. We were surprised by visual design changes across these papers, particularly the increasing prevalence of highly complex figures, and decided to investigate further by conducting semi-structured interviews with researchers who had been active in computer vision, graphics, and visualization for several decades, as well as computational analysis of a corpus of research papers published between 2013 and 2021 in the IEEE/CVF Conference on Computer Vision and Pattern Recognition—the highest-ranked publication venue in computer science and second highest in all of science, according to Google Scholar as of August 2024 (goo, 2024). We study this conference, rather than similar conferences in machine learning or natural language processing, because of the importance of visual presentation in computer vision. Along the way, our visual analysis, interviews, and computational findings coalesced into a media archaeology approach.
@@ 3.1. Media Archaeology
Inspired by recent work on the history of computer graphics (Gaboury, 2021), we examine historical research papers through the lens of media archaeology. “Archaeology” here refers to Foucault’s concept of archaeology as the “uncovering of the archive” (Foucault, 1969, p. 131), rather than physical excavation. Media archaeology, as defined by Huhtamo and Parikka, “rummages textual, visual and auditory archives as well as collections of artifacts, emphasizing both the discursive and material manifestations of culture” (Huhtamo and Parikka, 2011, p. 3). In other words, we study digital media objects as components of a media culture: a system of practices and meanings which structure our interpretation (Rose, 2016). We also consider their digital materiality as physical patterns of bits on hard drives, produced through the labor of researchers. By drawing attention to material culture, media archaeology is in dialogue with theories of historical materialism, which consider economic factors—production and consumption of goods and services—as the driving force of history, rather than revolutionary ideas or progress narratives (Gaboury, 2021, p. 5).
@@
In addition to our computational analysis and detailed study of the CVPR proceedings, we “rummaged” research papers from the IEEE, Computer Vision Foundation, and ACM digital archives, including the IEEE International Conference on Computer Vision (ICCV), European Conference on Computer Vision (ECCV), IEEE Transactions on Pattern Analysis and Machine Intelligence (PAMI), and the ACM Special Interest Group on Graphics and Interactive Techniques conference (SIGGRAPH). We examined ICCV, ECCV, and PAMI in addition to CVPR because they are the four most influential computer vision venues according to Google Scholar,111https://web.archive.org/web/20220709182958/https://scholar.google.com/citations?view_op=top_venues&hl=en&vq=eng_computervisionpatternrecognition and SIGGRAPH due to its historical significance and prevalence in our interviews. We turned to these venues to build context for the results we were seeing in CVPR. While we did not systematically code the enormous number of papers published in these venues, we collected dozens of screenshots of interesting figures and tables and discussed their visual style in weekly meetings, integrating interview and computational results which eventually coalesced into historical narratives.
@@ 3.2. Interviews
To augment our visual analysis, we conducted interviews with veteran computer vision researchers. As there are a relatively small number of eligible participants, we recruited individuals via email and in-person at the CVPR 2022 conference. Participants are listed by years of industry and academic experience in Table 1. This sample is highly non-representative, showing both survival bias based on who remains in the field over a long period of time and selection bias based on who agreed to an interview, so it does not serve as the sole source of our claims. Instead, it provides important social context to our visual analysis.
@@
After transcribing the interviews, we engaged in a process of iterative memoing, integrated into our weekly meetings where we compared examples from historical research papers and interview excerpts which shared common themes. We focused our analysis on identifying patterns from the context that participants provided while telling stories about their writing, as well how participants thought about their papers in relation to changing disciplinary practices, rather than simply analyzing the papers participants discussed or discussing our participants’ analysis of their own work.
@@ 3.3. Computational Analysis
As both qualitative approaches rely on the analysis of specific examples, we used supplementary quantitative analysis to verify that phenomena we observed in the interviews and media archaeology were actually as pervasive as participants seemed to think. Specifically, we were interested in whether teaser images, figures and tables were becoming more prevalent in CVPR papers over time, and whether more titles were following a particular format containing an acronym followed by a colon. To answer these questions, we collected PDFs of papers from the Computer Vision Foundation’s Open Access Repository (thecvf.com) published in CVPR 2013 to 2021 and article metadata from IEEE Xplore (ieeexplore.ieee.org) for CVPR 1992 to 2020. We manually inspected PDFs from thecvf.com to count teaser images, then used the Linux pdftotext tool combined with regular expressions to count figures and tables. From the Xplore metadata, we used regular expressions to parse paper titles, treating a word with more than two capital letters as an acronym, and an acronym as unique if it only appears once in a given year of data.
@@ 4. Findings
Through our analysis, we found that several elements of the contemporary computer vision research paper serve as material traces of the disciplinary change which took place during the 2010s. We find that papers are written to promote a specific contribution which is increasingly commodified. By “contribution,” we mean the model, algorithm, method, dataset, or other system that the paper offers to readers. By commodified, we mean reduced to its exchange value in terms of the attention it generates. As Simon theorized (Simon, 1971), and authors including Goldhaber (Goldhaber, 1997) and Davenport and Beck (Davenport and Beck, 2001) have applied to the web, when “capital, labor, information and knowledge are all in plentiful supply” the limiting factor is the attention of consumers, which begins to be treated like a commodity (Davenport and Beck, 2001).
However, this concept of the attention economy presupposes that attention is valuable, which is unhelpful for understanding why academics seek the attention of their peers instead of the larger research community or general public. To better understand why academic attention is uniquely valuable, we employ Bueno’s explanation, that online attention, in addition to a commodity, is a form of labor (Bueno, 2017). This claim relies on Romano Alquati’s concept of “valorization information” via Pasquinelli (Pasquinelli, 2015). Alquati, in his study of an Olivetti computer factory, found that workers contribute to mostly-automated production processes through their micro-decisions, which served as a feedback-control signal for the production process. Bueno argues that attention can provide a similar signal. Specifically, by monitoring attention, companies can forecast consumer behavior and more efficiently follow market trends. By doing so, these companies extract surplus value from online behavior data (Bueno, 2017, Ch. 2). We believe a similar process may be at play in academic publishing, where authors can leverage highly specialized early attention to increase the prestige value of their research. This process can take the form of feedback, where authors improve their work based on an early online response, or simply where early viewers find a paper interesting and share it, judging it as valuable, putting it in front of more eyes and increasing the chance that it will reach the right audience. Eventually, that attention translates into citations, jobs, students, grants, and other benefits of academic prestige.
We study three aspects of contemporary research papers: the teaser image, the results table, and the high resolution figure. We have grouped our results into thematic sections surrounding these three elements in relation to the concept of a paper’s contribution.
@@ 4.1.1. Teaser images
The “teaser image” is a large first-page figure which summarizes the paper. Several examples of teaser images are shown in Figures 1 and 2. These figures are functionally similar to the trend of “visual abstracts” (Ramos and Concepcion, 2020) and “table of contents images” (Buriak, 2011) in the biomedical sciences; however unlike these forms which are graphical summaries of the text, teaser images are part of the main paper and are usually a visualized system output that the authors want to show. Teaser images have been steadily gaining popularity at CVPR, as shown in Figure 3 (left). For this analysis, we define the teaser image as a first-page figure which covers the entire width of the page, not just a column.
@@
P12 ascribes the teaser image to the famous graphics researcher Randy Pausch: “Randy Pausch…takes credit for putting the teaser image in SIGGRAPH papers…he claimed that he did a paper where half the first page was a teaser image and after that that became the norm where people started always putting these images.” The paper in question is pictured in Figure 2 (a). While Pausch did not invent the large first-page figure (Tim Berners-Lee famously used one in his 1989 proposal for the world wide web (Berners-Lee, 1989)), he did publish the first paper in SIGGRAPH with this layout, and it quickly became popular. P2 and P10 point to the quick adoption of teaser images, which even became institutionalized in the SIGGRAPH template and later the template for all ACM conferences (Figure 2 (c)).
@@
P12 defines the teaser as a visualization of either a result or a system and explains that it has spread to many other conferences because of the way it attracts attention: “It’s a trailer. It’s to get people in…I think it’s a very compelling way to convey what the paper does.” For her, the teaser is a highly effective innovation which improves research papers. P2 echoes that sentiment: “They just made the papers look good! I mean, it’s much more memorable and there are some papers still today that I don’t remember the title, but you see the picture and you’re like, ‘Oh yeah, that’s the Randy Pausch paper on the VR for whatever, right?’ So yeah there’s a few of those that are just like, really iconic first pagers.” P2 is referencing the same Pausch paper as P12 (Pausch et al., 1996), highlighting its memorability.
@@
The teaser image is a trailer, a hook; it advertises the paper to the potential reader. The authors want to promote their paper and showcase the best results they can because the sheer number and easy access to papers has made it harder to stand out. The visual organization of these figures echoes that theme. Notice the commonalities between the two teaser images in Figure 2 (d) and (e). Both show the output of several algorithms attempting to solve the same problem, but they depict the results in such a way that their method is clearly best. This visual effect mirrors that of advertisements, which depict two competing products in action (Figure 2 (f)). While the experiments in research papers are far more sophisticated than those in advertisements, they share a common visual design. Along these lines, we find a recurring theme of teaser images which rely on the iconography of the 3D-reconstructed, brightly colored human body as a particularly compelling visual. Several examples are shown in Figure 1. Human faces and bodies have been found to attract consumer attention to advertisements (Wilkinson and Light, 2011; Ohme et al., 2011); a similar principle may be motivating computer vision authors.
@@
We observe a similar advertising quality in titles. Figure 3 (right) shows the rise in popularity of a particular title construction where an acronym, which is usually the name of a model, is followed by a colon: “HOPE-Net: A Graph-Based Model for Hand-Object Pose Estimation” (Doosti et al., 2020) or “DeMoN: Depth and Motion Network for Learning Monocular Stereo” (Ummenhofer et al., 2017). These names both signal that the paper proposes a new model and brands the paper with a short, memorable name which is often a cultural reference or clever pun. An especially cheeky example is Joseph Redmon and Ali Farhadi’s 2017 paper “YOLO9000: Better Faster Stronger” (Redmon and Farhadi, 2017), which improves upon an earlier model called YOLO (short for You Only Look Once) (Redmon et al., 2016), and references a variety of memes from the mid-2010s.222YOLO itself was a meme, short for “you only live once,” the number 9000 references a famous quote from the television show Dragon Ball Z and the phrase “better, faster, stronger” references a Daft Punk song. While unusual in computer vision, this kind of nerdy referential humor has been observed in other areas, for example, web design (Goree et al., 2022).